OpenStudy (anonymous):

Please help :O Given: sinA=-3/5, cosB=15/17, angle A in Quadrant III and angle B in Quadrant I. tan(A-B)= a. -13/36 b. -77/36 c. 13/84 d. 77/84

OpenStudy (anonymous):

b

OpenStudy (anonymous):

Thanks! Can you please show me how you got it?

OpenStudy (anonymous):

sin^1(-3/5)=-36.87=A cos^1(15/17)=28.07=B tan(A-B)=Tan(-36.87-28.07) =-2.138 =-77/36

jimthompson5910 (jim_thompson5910):

or use the idea that \[\Large \tan(A-B) = \frac{\tan(A)-\tan(B)}{1+\tan(A)\tan(B)}\]

jimthompson5910 (jim_thompson5910):

Then you can use the identity \[\Large \tan(x) = \frac{\sin(x)}{\cos(x)}\]

OpenStudy (anonymous):

or you could do wat jim said

OpenStudy (anonymous):

THANKS BOTH!

jimthompson5910 (jim_thompson5910):

np